If the relative errors of three quantities to be multiplied together are 0.001, 0.008 and 0.007, what is the relative error of the resulting quantity? Give your answer to three decimal places Answer:
Answer #1

Let the quantity be A which is given by

A = p\times q\times r ............(1)

Let the absolute errors in measurement of A, p q, and r be \Delta A, \Delta p , \Delta q, \Delta r respectively.

hence,

{A}\pm \Delta A = (p\pm \Delta p) \times (q\pm\Delta q) \times (r\pm \Delta r) \\ \Rightarrow \cancel{A}\pm \Delta A = \cancel{pqr} \pm p \Delta q r \pm p\Delta q \Delta r \pm pq\Delta r \pm \Delta p qr\pm \Delta p \Delta q r \pm \Delta p \Delta q \Delta r \pm \Delta p q \Delta r \\ \Rightarrow \pm \Delta A = \pm p \Delta q r \pm p\Delta q \Delta r \pm pq\Delta r \pm \Delta p qr\pm \Delta p \Delta q r \pm \Delta p \Delta q \Delta r \pm \Delta p q \Delta r ........(2)

Dividing equation (2) by equation (1), we get

\pm \frac{\Delta A}{A} = \pm \frac{\Delta q}{q} \pm \frac{\Delta q \Delta r}{qr} \pm \frac{\Delta r}{r} \pm \frac{\Delta p}{p} \pm \frac{\Delta p \Delta q}{pq} \pm \frac{\Delta p \Delta q \Delta r}{pqr} \pm \frac{\Delta p \Delta r}{pr} \\

Now, we can write this properly as

\pm \frac{\Delta A}{A} = \pm \frac{\Delta p}{p} \pm \frac{\Delta q}{q} \pm \frac{\Delta r}{r} \pm\left ( \frac{\Delta p}{p} \times \frac{\Delta q}{q} \right ) \pm \left ( \frac{\Delta q}{q} \times \frac{\Delta r}{r} \right ) \pm \left ( \frac{\Delta p}{p} \times \frac{\Delta r}{r} \right ) \pm \left ( \frac{\Delta p}{p} \times \frac{\Delta q}{q} \times \frac{\Delta r}{r} \right ) .......(3)

We know that the relative errors are given by

Relative \ error \ in \ A = \pm \frac{\Delta A}{A}

Relative \ error \ in \ p = \pm \frac{\Delta p}{p} = 0.001

Relative \ error \ in \ q = \pm \frac{\Delta q}{q} = 0.008

Relative \ error \ in \ r = \pm \frac{\Delta r}{r} = 0.007

hence, we can calculat the relative error in the measurement of the quantity using equation (3)

\frac{\Delta A}{A} = 0.001 + 0.008+ 0.007 +(0.001\times 0.008) + (0.008\times 0.007) +( 0.001\times 0.007) + (0.001\times 0.008\times 0.007)\\ \Rightarrow \frac{\Delta A}{A} = 0.016 + 0.000008 + 0.000056+ 0.000007+ 0.000000056 = 0.016071056 \\ \Rightarrow \frac{\Delta A}{A} \approx 0.016

Hence, the relative error in the measurement of the quantity is 0.016.

Note that, all other terms that appear in the relative error calculation dimish because of the product. Only the sum of individual relative errors of 3 quantities is good enough estimate the relative error of product.